A sample of chlorine gas has a volume of 6 liters at a temperature of 10 °C. What volume does the gas occupy at 50 °C?
1
Expert's answer
2020-10-26T14:43:54-0400

V1 = 6 L

T1 = 10°C = 283 K

V2 = unknown

T2 = 50°C = 323 K

P = const


Solution:

Since the pressure doesn't change, Charle's law can be used.

Charle's law law can be expressed as: V1/T1 = V2/T2

To find the final volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (6 L × 323 K) / (283 K) = 6.848 L = 6.85 L

V2 = 6.85 L


Answer: A sample of chlorine gas occupy 6.85 L at 50 °C.